# SAR Strikes Back: A New Hope for RSVQA

### v1

@@ Abstract
Remote sensing visual question answering (RSVQA) is a task that automatically extracts information from satellite images and processes a question to predict the answer from the images in textual form, helping with the interpretation of the image. While different methods have been proposed to extract information from optical images with different spectral bands and resolutions, no method has been proposed to answer questions from Synthetic Aperture Radar (SAR) images. SAR images capture electromagnetic information from the scene, and are less affected by atmospheric conditions, such as clouds. In this work, our objective is to introduce SAR in the RSVQA task, finding the best way to use this modality.
@@
We propose two different models to include the SAR modality. The first one is an end-to-end method in which we add an additional encoder for the SAR modality. In the second approach, we build on a two-stage framework. First, relevant information is extracted from SAR and, optionally, optical data. This information is then translated into natural language to be used in the second step which only relies on a language model to provide the answer.
We find that the second pipeline allows us to obtain good results with SAR images alone. We then try various types of fusion methods to use SAR and optical images together, finding that a fusion at the decision level achieves the best results on the proposed dataset. We show that SAR data offers additional information when fused with the optical modality, particularly for questions related to specific land cover classes, such as water areas.
keywords:
@@
journal: ISPRS Journal of Photogrammetry and Remote Sensing
[…]
{highlights}
We introduce the Synthetic Aperture Radar modality for Visual Question Answering
We propose two pipelines to take into account Synthetic Aperture Radar for the task
We propose a multi-modal setting adding optical imagery
We show that using both modalities allows to obtain better performances
1 Introduction
Public and private sector initiatives are facilitating access to a wide range of remote sensing images. A well-known example is the Sentinel satellite constellation launched in 2014 as part of the European Union’s Copernicus programme. The mission provides free access to a wide range of imagery, including optical and Synthetic Aperture Radar (SAR) images. The optical images acquired by Sentinel-2 deliver high-resolution information based on the reflected sunlight. In contrast, SAR images, captured with the Sentinel-1 satellites, use radar signals which penetrate clouds and darkness. They provide data on surface roughness, moisture content, and other physical properties through radar backscatter. Sentinel-1 images contain two polarization channels, VV and VH, offering a more comprehensive representation of the scene.
@@ 1 Introduction
However, a fusion of SAR with optical images can leverage the unique strengths of each type of imagery. SAR images provide complementary information, such as detailed texture and surface characteristics, which are often not visible in optical images due to atmospheric conditions or lighting. Integrating SAR with optical images can enhance overall analysis and interpretation by offering a more comprehensive view of the scene. This combined approach has been shown to improve various applications, including object detection and land classification  [10].
@@
The work is divided as follows: the state-of-the-art in RSVQA, deep learning with SAR images, and SAR/optical data fusion are presented in Section 2. In section 3, the two pipelines, End-to-End RSVQA and Prompt-RSVQA, and three fusion methods are presented, after we introduce our proposed dataset RSVQAxBEN-MM in section 4. Section 5 describes the scores used to evaluate the different steps of our methods and we present our experiments and the results in Section 6. These results are discussed in section 7, and finally in section 8 we propose our conclusions on the use of SAR in RSVQA and future perspectives in this field.
2 Related Works
@@
RSVQA has been first introduced in [5]. Since then, many methods have been proposed, including approaches for identifying the best language models [11], fusion method between text and images [12], using segmentation maps to guide the fusion [13], and using text-based data augmentation [14]. The applications field of RSVQA has been broadened, with models able to give different types of answers. For instance, in [15], an object detection model based on the textual question is examined. A model designed to answer questions requiring segmentation maps, vector maps, and object counting through a VQA interface, highlighting the capabilities of RSVQA is introduced in [16].
Research in RSVQA has been almost exclusively focused on optical images, with the exception of [8] in which SAR images are introduced. In recent years, efforts have been made to explore deep-learning based methods using SAR as an input. Hence, works on object detection [17], despeckling [18], images generation [19] and RSVQA with questions about the scattering patterns and the backscattering mechanisms present in the image  [20] have been proposed. The interaction between SAR images and text is only starting to be explored due to the inherent complexity of SAR data. However, deep learning seems to be ineffective at describing the size of targets or counting objects [21] using SAR images. On the other hand, it performs well in outlining the relative position of target objects, such as indicating whether an object is near others or providing density descriptions (e.g., ”few”, ”a lot” and ”many”) in SAR images. Text is also used with SAR to generate new images, to mitigate the imbalance in the dataset, and address the deficiency of insufficient data samples [22].
@@
One of the reasons for the predominant use of optical images is that they represent the visible spectrum. This makes them ideal for tasks like object recognition [23], though they are limited by weather and lighting conditions. In contrast, SAR images, generated using radar signals, provide structural information and operate effectively in all weather and lighting conditions but are harder to interpret due to their complex patterns. Combining the optical and SAR modalities can leverage the strengths of both modalities, with optical images providing visual information and SAR images offering structural insights, resulting in a more comprehensive and accurate representation of the scene [24].
@@
Various fusion techniques have been explored to combine optical and SAR images in deep learning pipelines. Early fusion, also known as raw data fusion, merges optical and SAR images before feature extraction, and has been applied in pansharpening [25] and Digital Surface Model (DSM) generation from multiangular images [26]. This approach is especially useful when the images are captured under similar conditions, such as using the same sensor or when the acquisition times are closely aligned. Halfway fusion, or feature-level fusion, extracts features from both modalities separately and then combines them for tasks like classification and change detection [27], offering a richer set of characteristics. Late fusion, or decision-level fusion, operates at the highest semantic level by merging the outputs of single-modality pipelines and is particularly effective in classification tasks [28]. The choice between halfway and late fusion, however, remains task-dependent. In classification, it is challenging to determine which technique performs better, as studies report that halfway fusion sometimes yields superior results [29, 30], while late fusion excels in other cases [31]. As noted in [32], the performance of either method is closely tied to the specifics of the problem and dataset. In [33], the authors demonstrate that a combination of all three fusion types can be advantageous, depending on the application.
It appears that there is a gap in the state-of-the-art regarding the use of SAR images for RSVQA, even though SAR has been demonstrated to be useful in several fields due to the complementary information it provides compared to optical images [24]. We propose to extend the work done in [8] by studying the behaviour of two pipelines when using SAR images alone or in addition to optical data.
3 Method
@@
The RSVQA task aims at providing an answer $a$ to a question $q$ from an image $i$. As seen in Section 2, recent works on RSVQA only focused on the case in which $i=i_{o}$ is an optical image (with multiple spectral channels $n_{o}$). In this work, we introduce a SAR image $i_{s}$ (with multiple polarization channels $n_{s}$) to the problem setting. Hence, we formulate our task as providing an answer $a$ to a question $q$ from either a single modality (i.e. $i=i_{s}$ or $i=i_{o}$) or a pair of optical and SAR images $i=(i_{o},i_{s})$. Similarly to other works, we frame the RSVQA task as a classification problem, where the answer $a$ is predicted among a set of pre-defined answers $\mathcal{A}$.
@@
To predict the right answer, we propose two methods shown in Figure 1: the first one is an end-to-end method (End-to-End RSVQA, described in Sub-section 3.1). In this method, abstract features are extracted from the images $i$, as well as from the question $q$. These features are then merged to predict the answer $a$. The second one (Prompt-RSVQA, described in Sub-section 3.2) is inspired by [7]. In this method, we first extract semantic information from the images. This is done through a separately trained multi-label classification network. This semantic information is then passed, along with the question $q$ to a language model to predict $a$. In this work, we also experiment with different methods for the extraction of the visual feature vector, with the proposition of three fusion mechanisms. The visual feature extractors are presented in Sub-section 3.3.
3.1 End-to-End RSVQA
@@
In the End-to-End RSVQA pipeline, both the images $i$ and the question $q$ are processed separately. The different encoders used to obtain $f_{i}$, the feature vector of the images $i$, are described in section 3.3. To obtain $f_{q}$, the feature vector of the question, we use a Recurrent Neural Network (RNN). We add a fully connected layer to map both vectors to new vectors $f^{\prime}_{i}$ and $f^{\prime}_{q}$ both of dimension $n_{a}$. Both of these vectors are merged into a new vector, $f_{a}$ through a point-wise multiplication. While the point-wise multiplication is a simple and fixed operation, the fully connected layers applied on $f_{i}$ and $f_{q}$ leave a degree of freedom for the network to reorganize the information in a way that helps the fusion of both vectors. Finally, the vector $f_{a}$, representing both the visual and textual information, is used as an input to a multi-layer perceptron to predict the final answer $a$. This MLP outputs a vector $y$, of size $|\mathcal{A}|$, giving a score for each possible answer from $\mathcal{A}$.
3.2 Prompt-RSVQA
@@
Prompt-RSVQA is organized into two stages. The first stage aims at extracting relevant semantic information from the image $i$. After obtaining $f_{i}$ (see Sub-section 3.3), we predict $f_{c}$ a vector that represents classification scores through a multi-layer perceptron. More specifically, after applying a threshold on the prediction vector $f_{c}$, we obtain the set of classes describing the content of the image, which are concatenated in a text-based list called context $c$. Because this operation is non-differentiable, this first part of the pipeline is trained separately.
@@ 3.3 Visual models
The second fusion method, halfway fusion (Figure 2(d)), involves processing the two modalities separately through two different CNNs. Before the final step, the two resulting feature maps are concatenated to form a single feature map $f_{i}$. Once again, the resulting feature vector is passed to a MLP to obtain $f_{c}$.
@@
4 Data
This section introduces a new dataset for multi-modal RSVQA: RSVQAxBEN-MM. This dataset is derived from three other datasets: BEN and BEN-MM, discussed in Sub-section 4.1 and RSVQAxBEN, discussed in Sub-section 4.2. A sample of each of these datasets is shown in Figure 3.
Figure 3: Workflow from the initial BEN dataset containing the Sentinel-2 optical images and their CLC classes, to BEN-MM where the respective Sentinel-1 images were added. We then move on to RSVQAxBEN in which questions and answers related to the optical images have been proposed. Finally, the dataset we propose in which such questions and answers are linked to both optical and SAR images.
4.1 Land-cover datasets: BEN and BEN-MM
@@
An extension of BEN, called BigEarthNet-Multi Modality dataset [35] (BEN-MM), has been later released adding to each Sentinel-2 patch of the original dataset the corresponding Sentinel-1 image, in the two polarisation channels (VV and VH) in dB. Note that this extension does not include new classes, and keeps the same 19 classes nomenclature as BEN. Moreover, the same train, validation and test sets distribution is kept. In this work, we also use another version of the BEN-MM dataset, with 61 classes as labels (see Section 4.2). In the rest of this manuscript, we refer to this dataset as BEN-MM-61.
4.2 RSVQA dataset: RSVQAxBEN
@@
The RSVQAxBEN dataset is derived from BEN. To each of the 590’326 Sentinel-2 RGB patches, it adds 25 question/answer pairs, for a total of 14’758’150 image/question/answer triplets. The questions are based on each image’s CLC labels. As opposed to BEN, the 64 original classes of the CLC nomenclature are kept. However, two classes with the same name at different hierarchy levels (water bodies and pastures) are counted as a single label and the category Glaciers and perpetual snow is removed, leading to a total of 61 classes.
@@
The questions can be divided into two types: yes/no questions, in which the answer is ’yes’ or ’no’ (e.g. ’Is there a forest or water in this image?’) and land cover questions in which the answer is a list of classes (e.g. ’Besides forest, what classes are present in the image?’). The dataset is heavily unbalanced, with 80.7% of the questions being a yes/no question. In yes/no questions, conjunctions such as ’and’ and ’or’, adding a difficulty level to the question, are present in 72.3% of the cases. In 27.1% of the questions, two conjunctions are found. As opposed to BEN, the dataset split is divided according to the spatial location of the image: The westernmost patches are placed in the training set (66% of the dataset); the easternmost patches are put in the test set (23% of the dataset); finally, the other patches, representing 11% of the dataset, are put in the validation set. This choice allows the models to avoid biases due to geographical location. However, it implies a difficult geographical generalization problem, as some classes present different visual representations in Western Europe (i.e. the train set) and in Eastern Europe (i.e. the test set).
4.3 Proposed dataset: RSVQAxBEN-MM
@@
This work introduces the RSVQAxBEN-MM dataset. This dataset adds the SAR images from BEN-MM to the question/answer of RSVQAxBEN. An example is shown in Figure 3(d). Because it is derived from RSVQAxBEN, this dataset keeps the same choices, in terms of classes of interest and train / validation / test sets distribution. Each of the images / question / answer triplets is composed of the Sentinel-2 RGB patch as in RSVQAxBEN and adds the corresponding Sentinel-1 patch. The provided Sentinel-1 patches are composed of the VV, and VH channels in dB and the ratio between the two, providing $n_{s}=3$ SAR channels. The ratio has proven to help classify the volume from the surface [8, 36]. Using images in dB, our ratio is a subtraction of the two channels normalized between 0 and 1.
4.4 Dataset Evaluation

### v2

@@ Abstract
Remote Sensing Visual Question Answering (RSVQA) is a task that automatically extracts information from satellite images. It then processes a question to predict the answer from the images in textual form, helping with the interpretation of the image.
While different methods have been proposed to extract information from optical images with different spectral bands and resolutions, only recently have some preliminary studies started exploring very high-resolution Synthetic Aperture Radar (SAR) data. These studies leverage SAR’s ability to capture electromagnetic information and operate in all atmospheric conditions. However, no research has compared the results obtained using SAR and optical imagery or explored methods to fuse the two modalities effectively.
This work investigates the integration of SAR images into the RSVQA task exploring the most effective way to combine them with optical images.
@@
We study two different pipelines for RSVQA to include SAR modality and introduce a dataset enabling SAR-based RSVQA. The first model is an End-to-End approach while the second is a two-stage framework. In the latter, relevant information is extracted from SAR, before being translated into natural language to be used in the second step which only relies on a language model to provide the answer.
Our results show that the second pipeline achieves strong performance using SAR alone, yielding an improvement of nearly 10% in overall accuracy compared to the first one. We then explore various types of fusion methods to use SAR and optical images together. A fusion at the decision level achieves the best results on the proposed dataset, with a final F1-micro score of 75.00% and an F1-average of 81.21% for classification, as well as an overall accuracy of 75.49% for VQA. We show that SAR data offers additional information when fused with the optical modality, particularly for questions related to specific land cover classes, such as water areas.
Index Terms:
@@
I Introduction
Public and private sector initiatives are facilitating access to a wide range of remote sensing images. A well-known example is the Sentinel satellite constellation launched in 2014 as part of the European Union’s Copernicus programme. The mission provides free access to a wide range of imagery, including optical and Synthetic Aperture Radar (SAR) images. The optical images acquired by Sentinel-2 deliver high-resolution information based on the reflected sunlight. In contrast, SAR images, captured with the Sentinel-1 satellites, use radar signals which penetrate clouds and operate effectively at night. They provide data on surface roughness, moisture content, and other physical properties through radar backscatter. Sentinel-1 images contain two polarization channels, VV and VH, offering a more comprehensive representation of the scene.
@@ I Introduction
However, a fusion of SAR with optical images can leverage the unique strengths of each type of imagery. SAR images provide complementary information, such as detailed texture and surface characteristics. In addition, the content of optical images may not be visible at all due to atmospheric conditions or poor lighting. Integrating SAR with optical images can enhance overall analysis and interpretation by offering a comprehensive view of the scene. This combined approach has been shown to improve various applications, including object detection and land classification  [10].
@@
The paper is structured as follows: Section II reviews RSVQA, deep learning with SAR, and data fusion. Section III presents our pipelines and fusion methods, followed by our dataset in Section IV. Section V details evaluation metrics, with experiments and results in Section VI. We discuss findings in Section VII and conclude in Section VIII with insights on SAR in RSVQA and future directions.
II Related Works
@@
RSVQA has been first introduced in [5]. Since then, many methods have been proposed, including approaches for identifying the best language models [11], fusion method between text and images [12], using segmentation maps to guide the fusion [13], and using text-based data augmentation [14]. The applications of RSVQA have broadened, with models now capable of providing diverse answers. For example, [15] explores an object detection model based on textual questions, while [16] introduces a model that answers questions requiring segmentation maps, vector maps, and object counting through a VQA interface, showcasing the versatility of RSVQA.
Research in RSVQA has predominantly focused on optical images, with SAR images being introduced only recently. The first study to incorporate SAR images [8], addressed land classification. Subsequent works explored questions related to scattering patterns and backscattering mechanisms [17], boat detection [18], and comparisons between the results got using optical images, obscured by clouds or captured at night, and SAR’s one [19].
Recent efforts in the remote sensing community have explored deep learning with SAR data. Hence, works on object detection [20], despeckling [21], volcano deformation detection [22] and images generation [23] have been proposed. In general, all techniques exclusively applied to optical imaging are now also extended to SAR imaging [9].
The interaction between SAR images and text is still in its early stages, primarily due to the inherent complexity of SAR data. While deep learning has shown limited effectiveness in tasks such as describing target sizes or counting objects [24], it performs better when identifying the relative positions of targets—such as indicating proximity to other objects—or providing density-based descriptions like ”few,” ”many,” or ”a lot.” Additionally, text has been used in conjunction with SAR data to generate synthetic images, helping to mitigate dataset imbalances and address the challenge of limited training samples [25].
@@
One of the reasons for the predominant use of optical images is that they represent the visible spectrum. This makes them ideal for tasks such as object recognition [26], though they are limited by weather and lighting conditions. In contrast, SAR images, generated using radar signals, provide structural information and operate effectively in all weather and lighting conditions but are harder to interpret due to their complex patterns. Combining the optical and SAR modalities can leverage the strengths of both modalities, with optical images providing visual information and SAR images offering structural insights, resulting in a more comprehensive and accurate representation of the scene [27].
@@
Various fusion techniques have been explored to combine optical and SAR images in deep learning pipelines. Early fusion, also known as raw data fusion, merges optical and SAR images before feature extraction, and has been applied in pansharpening [28] and Digital Surface Model (DSM) generation from multiangular images [29]. This approach is especially useful when the images are captured under similar conditions, such as using the same sensor or when the acquisition times are closely aligned. Halfway fusion, or feature-level fusion, extracts features from both modalities separately and then combines them for tasks like classification and change detection [30], offering a richer set of characteristics. Late fusion, or decision-level fusion, operates at the highest semantic level by merging the outputs of single-modality pipelines and is particularly effective in classification tasks [31]. The choice between halfway and late fusion, however, remains task-dependent. In classification, it is challenging to determine which technique performs better, as studies report that halfway fusion sometimes yields superior results [32, 33], while late fusion excels in other cases [34]. As noted in [35], the performance of either method is closely tied to the specifics of the problem and dataset.
Although some recent works have explored RSVQA with SAR images, all focus on very high-resolution SAR images [17] [18] [19]. This makes it difficult to compare the results with those obtained from optical images captured in clear weather and daylight conditions, especially when the questions differ for SAR and optical images and the images do not represent the same areas [36].
While it is known that SAR images perform better at night and on cloudy days, there has been no study in the RSVQA field on whether they provide unique information that optical images do not. Our work aims to address this gap by building on the findings of [8], examining the performance of two RSVQA pipelines when using SAR images alone or in combination with optical data.
III Method
@@
The RSVQA task aims at providing an answer $a$ to a question $q$ from an image $i$. As seen in Section II, recent works on RSVQA only focused on the case in which $i=i_{o}$ is an optical image (with multiple spectral channels $n_{o}$). In this work, we introduce a SAR image $i_{s}$ (with multiple polarization channels $n_{s}$) to the problem setting. Hence, we formulate our task as providing an answer $a$ to a question $q$ from either a single modality (i.e. $i=i_{s}$ or $i=i_{o}$) or a pair of optical and SAR images $i=(i_{o},i_{s})$. Similarly to other works, we frame the RSVQA task as a classification problem, where the answer $a$ is predicted among a set of pre-defined answers $\mathcal{A}$.
@@
To predict the right answer, we propose two methods shown in Figure 1: the first one is an End-to-End method (End-to-End RSVQA, described in Sub-section III-A). In this method, abstract features are extracted from the images $i$, as well as from the question $q$. These features are then merged to predict the answer $a$. The second one (Prompt-RSVQA, described in Sub-section III-B) is inspired by [7]. In this method, we first extract semantic information from the images. This is done through a separately trained multi-label classification network. This semantic information is then passed, along with the question $q$ to a language model to predict $a$. In this work, we also experiment with different methods for the extraction of the visual feature vector, with the proposition of three fusion mechanisms. The visual feature extractors are presented in Sub-section III-C.
III-A End-to-End RSVQA
@@
In the End-to-End RSVQA pipeline, both the images $i$ and the question $q$ are processed separately. The different encoders used to obtain $f_{i}$, the feature vector of the images $i$, are described in section III-C. To obtain $f_{q}$, the feature vector of the question, we use a Recurrent Neural Network (RNN). We add a fully connected layer to map both vectors to new vectors $f^{\prime}_{i}$ and $f^{\prime}_{q}$ both of dimension $n_{a}$. Both of these vectors are merged into a new vector, $f_{a}$ through a point-wise multiplication. While the point-wise multiplication is a simple and fixed operation, the fully connected layers applied on $f_{i}$ and $f_{q}$ leave a degree of freedom for the network to reorganize the information in a way that helps the fusion of both vectors. Finally, the vector $f_{a}$, representing both the visual and textual information, is used as an input to a multi-layer perceptron to predict the final answer $a$. This MLP outputs a vector $y$, of size $|\mathcal{A}|$, giving a score for each possible answer from $\mathcal{A}$.
III-B Prompt-RSVQA
@@
Prompt-RSVQA is organized into two stages. The first stage aims at extracting relevant semantic information from the image $i$. After obtaining $f_{i}$ (see Sub-section III-C), we predict $f_{c}$ a vector that represents classification scores through a multi-layer perceptron. More specifically, after applying a threshold on the prediction vector $f_{c}$, we obtain the set of classes describing the content of the image, which are concatenated in a text-based list called context $c$. Because this operation is non-differentiable, this first part of the pipeline is trained separately.
@@ III-C Visual models
The second fusion method, halfway fusion (Figure 2(d)), processes the two modalities separately through different CNNs. Before the final step, their feature maps are concatenated into a single feature map $f_{i}$, then passed to an MLP to obtain $f_{c}$.
@@
IV Data
This section introduces a new dataset for multi-modal RSVQA: RSVQAxBEN-MM. This dataset is derived from three other datasets: BEN and BEN-MM, discussed in Sub-section IV-A and RSVQAxBEN, discussed in Sub-section IV-B. A sample of each of these datasets is shown in Figure 3.
Figure 3: Workflow from the initial BEN dataset containing the Sentinel-2 optical images and their CLC classes, to BEN-MM-61 where the respective Sentinel-1 images were added. We then move on to RSVQAxBEN in which questions and answers related to the optical images have been proposed. Finally, the dataset we propose in which such questions and answers are linked to both optical and SAR images.
IV-A Land-cover datasets: BEN and BEN-MM
@@
An extension of BEN, called BigEarthNet-Multi Modality dataset [38] (BEN-MM), has been later released adding to each Sentinel-2 patch of the original dataset the corresponding Sentinel-1 image, in the two polarisation channels (VV and VH) in dB. In this work, we also use another version of the BEN-MM dataset, with 61 classes as labels (see Section IV-B). In the rest of this manuscript, we refer to this dataset as BEN-MM-61.
IV-B RSVQA dataset: RSVQAxBEN
The RSVQAxBEN dataset is derived from BEN. Each of the 590,326 Sentinel-2 RGB patches is paired with 25 question/answer pairs, totaling 14,758,150 image/question/answer triplets.
@@ IV-B RSVQA dataset: RSVQAxBEN
The questions are based on each image’s CLC labels. As opposed to BEN, the 64 original classes of the CLC nomenclature are kept. However, two classes with the same name at different hierarchy levels (water bodies and pastures) are counted as a single label and the category Glaciers and perpetual snow is removed, leading to a total of 61 classes.
@@
Questions fall into two categories: yes/no questions, in which the answer is ’yes’ or ’no’ (e.g. ’Is there a forest or water in this image?’) and land cover questions in which the answer is a list of classes (e.g. ’Besides forest, what classes are present in the image?’). The dataset is heavily unbalanced, with 80.7% of the questions being a yes/no question. In yes/no questions, conjunctions such as ’and’ and ’or’, adding a difficulty level to the question, are present in 72.3% of the cases. In 27.1% of the questions, two conjunctions are found. As opposed to BEN, the dataset split is divided according to the spatial location of the image: The westernmost patches are placed in the training set (66% of the dataset); the easternmost patches are put in the test set (23% of the dataset); finally, the other patches, representing 11% of the dataset, are put in the validation set. This choice allows the models to avoid biases due to geographical location. This split minimizes location-based biases but introduces a geographical generalization challenge, as some classes vary visually between Western and Eastern Europe.
IV-C Proposed dataset: RSVQAxBEN-MM
@@
This work introduces the RSVQAxBEN-MM dataset. This dataset adds the SAR images from BEN-MM to the question/answer of RSVQAxBEN. An example is shown in Figure 3(d). Because it is derived from RSVQAxBEN, this dataset keeps the same choices, in terms of classes of interest and train / validation / test sets distribution. Each of the images / question / answer triplets is composed of the Sentinel-2 RGB patch as in RSVQAxBEN and adds the corresponding Sentinel-1 patch. The provided Sentinel-1 patches are composed of the VV, and VH channels in dB and the ratio between the two, providing $n_{s}=3$ SAR channels.
We use three channels, as combining VV and VH polarizations improves the separability of vegetation and artificial structures [39], while the VV/VH ratio effectively distinguishes volume from surface scattering [8, 40]. Both enhancing land cover classification in RSVQA tasks. In this context, with images expressed in dB scale, the ratio channel is calculated as the difference between VV and VH, normalized between 0 and 1.
The RGB representation of the SAR image is obtained by assigning VV to the R channel, VH to the green channel and the ratio to the blue channel.
IV-D Dataset evaluation)